# A Critical Look At Tokenwise Reward-Guided Text Generation

### v1

@@ 4 Related Work
Language model alignment Simple fine-tuning and instruction tuning [37] are ways to align LLMs to labeled data. Recently, RLHF methods [7, 43, 17, 17, 23, 30] have provided a direct method to align LLMs to human preferences. The PPO algorithm has specially been popular and has shown promising results for a range of tasks [1, 2, 24]. However, training RL models is compute intensive and researchers have turned their attention to supervised fine-tuning methods that can learn directly from preference data. Liu et al. [21] turns the preference data into prompts with which they fine-tune the LLM. Dong et al. [10] uses the reward model to filter the training set to better fine-tune the model. DPO models [27, 28] the language model as a Bradley-Terry model and optimizes the RLHF objective without any need for RL. These training based methods, however, still fine-tune the base LLM which can be expensive as we scale.
Guided decoding There has been prior work in guided decoding using sequence-level [38, 35, 19, 16, 18, 13, 40] and token-level [8, 15, 15, 39, 5, 22] value functions. These token-level guided decoding algorithms are different from our work as they do not typically align language models using reward models or preference datasets. Deng and Raffel [9] use a reward model in the decoding process however, they use a cumulative squared loss function that is different from the RLHF framework. The closest work to our method is Khanov et al. [14], whose work is based on the Bradley-Terry model, but they use a reward model trained on full-sequences which we argue can lead to pitfalls (Theorem 1). Concurrent to our work, Zhao et al. [41] present a reward guided decoding method based on sequential Monte Carlo and show that it can approximate RLHF.
@@ 5 Experiments
We evaluate our proposed approach, which we call Partial Alignment as Reward-Guided Sampling (PARGS) [^1] This is in contrast to ARGS which considers full sequences and greedy decoding instead of sampling. on two language generation tasks: summarization and dialogue generation.
5.1 Experimental Setup
Summarization task We use the Reddit TL;DR dataset [36], where, the context $x$ is a post on the Reddit forum and $y$ is the summary of the post. We use the human preference dataset from [31] to train the reward model and the relevant baselines. Our base summarization model is a GPT2222-large model fine-tuned on the TL;DR training set. We use a pretrained reward model based on the DeBerta-v3-large architecture [^2] OpenAssistant/reward-model-deberta-v3-large-v2 and train it with partial sequences for an additional epoch. Our baselines include top-$k$ sampling [11], Best-of-$N$ generation, which involves sampling $N$ sequences from reference LLM and returning the best one according to the reward model, RLHF models based on PPO [29] and DPO [27], and the reward-base decoding method ARGS [14].
Dialogue task Next we evaluate our model on single turn dialogue using the Anthropic Helpful and Harmless (HH) [2] dataset. The goal is to generate a helpful and harmless response to a general purpose query. Each sample provides a prompt ${\mathbf{x}}$ and two responses ${\mathbf{y}}$ with a label indicating the preferred response. Here ${\mathbf{x}}$ is the human dialogue and ${\mathbf{y}}$ is the response from the assistant. We use Llama-2-7b (7 billion) as the base model and DeBerta-v3 as the reward model which is about 20×20$\times$20 × smaller. Training details, including hyper-parameters are presented in Appendix A.
Evaluation Following [14] we compare the algorithms based on average reward on the test samples as measured by the reward model. A higher reward indicates better alignment with the reward model. Note that we use a different full-sequence reward model and not the partial-sequence reward model (that we trained for our algorithm) to evaluate the models. Moreover, evaluating language generation, especially unconditionally, is nuanced and human evaluation is generally preferred, but time consuming. An alternative is LLM based evaluation, which has been shown to align with human assessment [42, 27]. We adopt GPT-4 based evaluation as a proxy of human evaluation. Following [6] we construct prompts for the two tasks and ask GPT-4 to score and rank response pairs. We randomly shuffle the order of the responses to mitigate position bias [42]. Finally we use the Rouge-L [20] score to measure the diversity of our generations for the dialogue task.
5.2 Results
Table 1 (left) shows the average reward for the summaries generated by the different algorithms as measured by the reward model. PARGS achieves the best average reward among the techniques that keep the LLM frozen and generate a single response ${\mathbf{y}}$. For reference, we also note that PARGS is competitive with DPO and PPO based RLHF that incur a large cost to finetune the LLM and Best-of-N that incur significant overhead to generate multiple responses. Note that we also evaluate our algorithm with greedy decoding (PARGS-G) for a direct comparison with ARGS. Similarly, Table 1 (right) presents average rewards for the responses of the different algorithms on the HH dialogue task. Note that in this setting the reward model is 20×20$\times$20 × smaller than the base LLM. Again, PARGS achieved the highest reward among the techniques that freeze the LLM and generate a single response. We observe that Best-of-$N$ achieved the highest average reward followed by DPO, but incurred overhead to generate multiple responses and fine-tune the LLM respectively.
@@ 5.2 Results
Next we evaluate PARGS against Best-of-N, ARGS, DPO and PPO using GPT-4. The prompt used to probe GPT-4 is presented in Appendix D. Table 2 reports the win-tie rate (i.e., percentage of utterances where GPT-4 finds PARGS’ response to be better than or equivalent to those of the baselines). Table 2 (left) shows that PARGS has a higher win-tie rate compared to all the methods, especially ARGS, for RL;DR summarization. As noted by others [27], Best-of-$N$ is a strong baseline, but it is computationally intensive (inference cost is 2×2$\times$2 × that of PARGS as described in Section 5.3). On HH, we observe (Table 2 right) that PARGS is better than top-$k$ and ARGS, but worse than Best-of-$N$ and DPO. As we scale training based alignment methods, e.g., DPO become prohibitive. DPO and Best-of-$N$ also trade diversity for accuracy as discussed next.
@@
Top-$k$ 0.1946
DPO 0.2068
PARGS 0.1881
Table 3: Diversity based on ROUGE-L.
We evaluate the diversity of generation on HH dialogue generation. We compare the sampling based techniques by generating $10$ responses for each prompt, evaluating the Rouge-L score between each generated pair and averaging the score. A lower Rouge-L score indicates a higher diversity. We note in Table 3 that PARGS generates the most diverse responses compared to top-$K$ and DPO. Note that Best-of-$N$ generates $N\times$ the number of samples from top-$K$.
@@ 5.3 Decoding costs
We present an estimate for the floating point operations (FLOPs) per token for inference with PARGS. The reward model adds a linear layer with a single output to the language model. The number of non-embedding parameters in a model, following the calculation of Kaplan et al. [12], is approximately $N\approx 12n_{\text{layers}}d_{\text{model}}^{2}$, where $n_{\text{layers}}$ is the number of layers and $d_{\text{model}}$ is the hidden dimension size. Additionally the FLOPs required by a forward pass is $C_{\text{forward}}\approx 2N+2n_{\text{layers}}n_{\text{ctx}}d_{\text{model}}$, where $n_{\text{ctx}}$ is the number of context tokens. The additional operations include $4d_{\text{model}}$ for the embedding and $2d$ for reward predicting. But since $6d_{\text{model}}\ll N$, $C_{\text{RM}}\approx C_{\text{forward}}$. Also if $d_{\text{model}}\gg\nicefrac{{n_{\text{ctx}}}}{{12}}$ we can assume that $C_{\text{RM}}=C_{\text{forward}}=2N$ [9]. At decode time we analyse $k$-tokens using the reward model. In our experiments $k=10$, so the total inference cost is $C_{\text{forward}}+10C_{\text{RM}}$ FLOPs per token. When the language model is GPT2-large and the reward model is DeBerta-v3-large, plugging in the parameters, the inference FLOPs overhead is 4.3×4.3$\times$4.3 × the base model. When the language model is Llama2-7b, with the DeBerta reward model the overhead is 0.47×0.47$\times$0.47 ×. Note that the Best-of-$N$ decoding cost overhead would always be 9×9$\times$9 ×. Appendix B shows the runtime to generate each token.
@@ 5.4 Limitations
Ideally, human evaluators are used to judge the generated sequences. However, we did not do so due to the monetary costs of conducting such an experiment. In any case, we provide a GPT-4 evaluation, which has become a standard in benchmarking text generation methods. Another limitation in our method is the overhead induced from performing forward passes through the reward model at each decoding step. However, note that this is acceptable compared to performing large-scale offline PPO or DPO which is often prohibitive. Moreover, this limitation is shared with other RGTG methods.
@@ 6 Conclusion
We have discussed the pitfalls in tokenwise, decoding-time reward-guided text generation (RGTG) with reward models trained on full sequences. These pitfalls can lead to inadequate reward during the autoregressive decoding process and may lead to subpar performance. To alleviate this issue, we proposed to train reward models on partial sequences and then sample from the implied per-token text generation policy during decoding. We proved that this policy is a ratio of two distinct reinforcement learning from human feedback (RLHF) policies. This means that this policy is not equivalent to the standard offline RLHF methods. However, we have also shown that it is intractable to obtain a tokenwise policy that is equivalent to a single RLHF policy. Training a partial-sequence reward model can thus be seen as a tradeoff between avoiding the pitfall of using full-sequence reward model in RGTG and tractability. Our experiment results validated our approach: it performs better than a recent RGTG method, ARGS, that leverages full-sequence reward models.
Acknowledgments and Disclosure of Funding
@@
Resources used in this work were provided by the Province of Ontario, the Government of Canada through CIFAR, companies sponsoring the Vector Institute https://vectorinstitute.ai/partners/ and the Natural Sciences and Engineering Council of Canada. AK thanks Rob Brekelmans for a fruitful discussion. AR thanks Apple for support through the Waterloo Apple PhD Fellowship, Natural Sciences and Engineering Council of Canada for its support through the PGS-D program, and the David R. Cheriton Graduate Scholarship. JG thanks Microsoft Research for support through its PhD Scholarship Programme and the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
Appendix A Training Details

### v2

@@ 4 Related Work
Language model alignment Simple fine-tuning and instruction tuning (Wei et al., 2021) are ways to align LLMs to labeled data. Recently, RLHF methods (Christiano et al., 2017; Ziegler et al., 2019a; Lee et al., 2021; Nakano et al., 2021; Snell et al., 2022) have provided a direct method to align LLMs to human preferences. The PPO algorithm has specially been popular and has shown promising results for a range of tasks (Askell et al., 2021; Bai et al., 2022; Ouyang et al., 2022). However, training RL models is compute intensive and researchers have turned their attention to supervised fine-tuning methods that can learn directly from preference data. Liu et al. (2023a) turns the preference data into prompts with which they fine-tune the LLM. Dong et al. (2023) uses the reward model to filter the training set to better fine-tune the model. DPO (Rafailov et al., 2023; 2024) models the LLM itself as a Bradley-Terry model and optimizes the RLHF objective without any need for RL. TDPO (Zeng et al., 2024) incorporates token-level KL divergence into the DPO objective to improve content diversity. These methods, however, still fine-tune the base LLM, which can be expensive as we scale. Some works have attempted to improve alignment by gathering more fine-grained rewards by using either LLMs (Cao et al., 2024) or human annotators (Wu et al., 2023).
Guided decoding There has been prior work in guided decoding using sequence-level (Welleck et al., 2022; Uesato et al., 2022; Lightman et al., 2023; Krishna et al., 2022; Li et al., 2023; Khalifa et al., 2023; Yao et al., 2023) and token-level value functions (Dathathri et al., 2019; Krause et al., 2021; Yang & Klein, 2021; Chaffin et al., 2022; Liu et al., 2023b). PPLM  (Dathathri et al., 2019) uses the gradients from an attribute classifier to guide LLM generation. Gedi (Krause et al., 2021) uses attribute conditioned language models as discriminators to update LLM generation probabilities using Bayes rule. These algorithms are different from our work as they do not align LLMs using human preference data. Deng & Raffel (2023) use a reward model trained on preference data in the decoding process, however they use a cumulative squared loss function that is different from the RLHF framework. Mudgal et al. (2024) uses a similar loss function with the key difference that instead of training with samples from a preference dataset, they take as input a full sequence reward model and train a partial sequence value function based on roll-outs (i.e., sampled token sequences) from the base LLM. Therefore, for each new base LLM, the value function needs to be retrained with new roll-outs, limiting portability to new or updated language models. The closest work to our method is Khanov et al. (2024), which is also based on the Bradley-Terry model, but they use a reward model trained on full sequences, which we have argued can lead to pitfalls. Different from our work, Zhao et al. (2024) a reward-guided decoding method based on sequential Monte Carlo and show that it can approximate RLHF.
Partial Rewards Outside of preference data alignment and RLHF, prior work in reinforcement learning for language modeling has looked at partial reward models for improving text generation. Hao et al. (2022) show that a sequence to sequence model trained with supervised learning is a valid partial reward model for text generation under a Markov decision process.  Lee et al. (2023) do not train an explicit reward model but instead introduce a ranking function which can rank the next token for partial sequences. Both these methods modify language model training.
@@ 5 Experiments
We evaluate our proposed approach, which we call Partial Alignment as Reward-Guided Sampling (PARGS)—in contrast to ARGS which considers full sequences and greedy decoding instead of sampling—on two language generation tasks: summarization and dialogue generation.
5.1 Setup
Summarization task We use the Reddit TL;DR dataset (Völske et al., 2017), where, the context ${\bm{x}}$ is a post on the Reddit forum and ${\bm{y}}$ is the summary of the post. We use the human preference dataset from Stiennon et al. (2020a) to train the reward model and the relevant baselines. Our base summarization model is GPT2-large, fine-tuned on the TL;DR training set. We use a pretrained reward model based on the DeBerta-v3-large architecture and train it with partial sequences for an additional epoch. Our baselines include top-$k$ sampling (Fan et al., 2018), Best-of-$N$ generation, which involves sampling $N$ sequences from reference LLM ($N=10$ for all our experiments) and returning the best one according to the reward model, RLHF models based on PPO (Schulman et al., 2017) and DPO (Rafailov et al., 2023), the reward-base decoding method ARGS (Khanov et al., 2024) and controlled decoding (CD; Mudgal et al., 2024). We use CD-Fudge as the baseline in all our CD experiments, noting that its performance is similar to CD-Q (see Table 4 in Mudgal et al. (2024)).
Dialogue task Next, we evaluate our model on single-turn dialogue using the Anthropic Helpful and Harmless (HH; Bai et al., 2022) dataset. The goal is to generate a helpful and harmless response to a general purpose query. Each sample provides a prompt ${\mathbf{x}}$ and two responses ${\mathbf{y}}$ with a label indicating the preferred response. We use Llama-2-7b as the base model and DeBerta-v3 as the reward model which is about 20×20$\times$20 × smaller. All details are in Appendix B.
Fine-grained text generation task We also evaluate our model on text generation using the UltraFeedback dataset (Ganqu Cui et al., 2024). We use Zephyr-7B as the base LLM and Phi-1.5 (1.3 billion parameters) as the reward model.
Machine translation task We perform additional experiments on machine-translation on the IWSLT-2017 dataset (Cettolo et al., 2017). We used the post-edit dataset from Kreutzer et al. (2020) on the IWSLT-2017 English-German dataset to provide token-wise reward signals. We use Gemma-2b as both the base model and the reward model. The evaluation is based on the standard BLUE score.
Evaluation Following Khanov et al. (2024), we compare all methods based on average reward on the test samples as measured by the reward model. We use a different full-sequence reward model and not the partial-sequence reward model (that we trained for our algorithm) to evaluate the models. Since evaluating language generation, especially unconditionally, is nuanced and human evaluation is very expensive, we use GPT-4-based evaluation, which has been shown to align with human assessment (Zheng et al., 2023; Rafailov et al., 2023). Following Chiang et al. (2023) we construct prompts for the two tasks and ask GPT-4 to score and rank response pairs. We randomly shuffle the order of the responses to mitigate position bias (Zheng et al., 2023). Finally, we use the Rouge-L score (Lin, 2004) and the BLEU score to evaluate the dialogue and translation tasks, respectively.
5.2 Results
Table 1 (left) shows the average reward for the summaries generated by the different algorithms as measured by the reward model. PARGS achieves the best average reward among the techniques that keep the LLM frozen and generate a single response ${\mathbf{y}}$. For reference, we also note that PARGS outperforms DPO and is competitive with PPO based RLHF that incurs a large cost to finetune the LLM, and Best-of-N that incurs significant overhead to generate multiple responses. Upon siginficance testing we observed PARGS to be significantly better than all algorithms expect PPO. Details in Appendix C. Note that we also evaluate our algorithm with greedy decoding (PARGS-G) for a direct comparison with ARGS.
Similarly, Table 1 (right) presents average rewards for the responses of the different algorithms on the HH dialogue task. Note that in this setting the reward model is 20×20$\times$20 × smaller than the base LLM. Again, PARGS achieved the highest reward among the techniques that freeze the LLM and generate a single response. We observe that Best-of-$N$ achieved the highest average reward followed by DPO, but incurred overhead to generate multiple responses and fine-tune the LLM respectively. Finally, Table 3 (left) presents average rewards on the UltraFeedback dataset. We observe that PARGS outperforms all methods except Best-of-$N$. Significance testing (see Appendix C) reveals that PARGS is significantly better.
Ultra Feedback
Method LLM Single ${\mathbf{y}}$? $r\pm\text{SE}$
Top-$k$ frozen yes -0.18$\pm$0.12
CD frozen yes -0.04$\pm$0.01
ARGS frozen yes 0.01$\pm$0.12
PARGS frozen yes 0.21$\pm$0.09
PARGS-G frozen yes 0.21$\pm$0.12
DPO trained yes -0.57$\pm$0.09
Best-of-$N$ frozen no 1.15 $\pm$0.08
Ultra Feedback
[…]
PARGS CD 53 - 13
PARGS ARGS 42 - 23
PARGS Best-of-$N$ 29 - 19
PARGS Top-$k$ 52 - 15
PARGS DPO 65 - 7
Table 3: Average reward (100 samples) $\pm$ std. error and GPT-4 evaluation for Ultra Feedback.
@@ 5.2 Results
Next we evaluate PARGS using GPT-4. The prompt used to probe GPT-4 is presented in Appendix H. Table 2 reports the win-tie rate (i.e., percentage of utterances where GPT-4 finds PARGS’ response to be better than or equivalent to those of the baselines). Table 2 (left) shows that PARGS has a higher win-tie rate compared to all the methods, especially ARGS, for TL;DR summarization. As noted by others Rafailov et al. (2023), Best-of-$N$ is a strong baseline, but it is computationally intensive. On HH, we observe (Table 2 right) that PARGS is better than CD and ARGS, but worse than Best-of-$N$ and DPO. As we scale training based alignment methods, e.g., DPO become prohibitive. On UltraFeedback (Table 3 right) we observe that PARGS outperforms all methods except Best-of-$N$. We perform human evaluation for PARGS against ARGS, CD and DPO on UltraFeedback. PARGS wins against all three baselines. Detailed results in Appendix E.
Table 4: BLEU Score on IWSLT-17 English to German
Method BLEU $\uparrow$
Greedy 31.7 $\pm$ 3.6
ARGS 29.4 $\pm$ 3.4
PARGS-G 33.2 $\pm$ 3.5
We next perform an experiment on English to German machine translation. We have access to a dataset of post-edits. The edited sequence is considered the winning sequence. Table 4 compares ARGS, PARGS-G and greedy decoding. We observe that applying ARGS reduces the BLUE score of the greedy baseline where PARGS-G increases it by 1.5 on average.
Table 5: Diversity based on ROUGE-L.
@@
Top-$k$ 0.230 ±$\pm 0.011$
DPO 0.206 ±$\pm 0.006$
PARGS 0.203 ±$\pm 0.008$
We evaluate the diversity of generation on 50 samples from the UltraFeedback dataset. We compare the sampling based techniques by generating $10$ responses for each prompt, evaluating the Rouge-L score between each generated pair. A lower Rouge-L score indicates a higher diversity. Table 5 shows that PARGS generates the most diverse responses compared to top-$K$ and DPO.
@@ 5.3 Decoding costs
We present an estimate for the floating point operations (FLOPs) per token for inference with PARGS. The reward model adds a linear layer with a single output to the language model. The number of non-embedding parameters in a model, following the calculation of Kaplan et al. (2020), is approximately $N\approx 12n_{\text{layers}}d_{\text{model}}^{2}$, where $n_{\text{layers}}$ is the number of layers and $d_{\text{model}}$ is the hidden dimension size. Additionally the FLOPs required by a forward pass is $C_{\text{forward}}\approx 2N+2n_{\text{layers}}n_{\text{ctx}}d_{\text{model}}$, where $n_{\text{ctx}}$ is the number of context tokens. The additional operations include $4d_{\text{model}}$ for the embedding and $2d$ for reward predicting. But since $6d_{\text{model}}\ll N$, $C_{\text{RM}}\approx C_{\text{forward}}$. Also if $d_{\text{model}}\gg\nicefrac{{n_{\text{ctx}}}}{{12}}$ we can assume that $C_{\text{RM}}=C_{\text{forward}}=2N$ (Deng & Raffel, 2023). At decode time we analyse $k$-tokens using the reward model. In our experiments $k=10$, so the total inference cost is $C_{\text{forward}}+10C_{\text{RM}}$ FLOPs per token. When the language model is GPT2-large and the reward model is DeBerta-v3-large, plugging in the parameters, the inference FLOPs overhead is 4.3×4.3$\times$4.3 × the base model. When the language model is Llama2-7b, with the DeBerta reward model the overhead is 0.47×0.47$\times$0.47 ×. Note that the Best-of-$N$ decoding cost overhead would always be 9×9$\times$9 ×. Appendix D shows the runtime to generate each token.
5.4 Partial Sequence Data Generation
We randomly sample a subset of the set of all partial sequences to maintain a reasonable computational budget. We present an ablation on the tldr summarization dataset where we present the average reward achieved by PARGS when training on different subsets, as well as the wall clock time.
Dataset Size Average Reward Wall clock time (approx)
1x 1.64 $\pm$ 0.22 1 hour
1.5x 2.32 $\pm$ 0.19 1.5 hour
2x 2.36 $\pm$ 0.20 2. hour
3x 2.23 $\pm$ 0.20 3 hour
Table 6: Average Reward with different partial sequence dataset sizes. x is the size of the full sequence dataset.
We can observe from the results on Table 6 that we get diminishing returns when the dataset is more than 1.5x. On the TLDR dataset we sample 2x and on the other datasets 1.5x of the total dataset size. Note that the wall clock time is for training on 4 RTX6000 GPUs.
@@ 5.5 Limitations
Ideally, all our experiments should be evaluated using human evaluation. However, we did not do so due to the monetary costs of conducting such an experiment and instead provide human evaluation for experiments on one of the dataset. In any case, we provide a GPT-4 evaluation, which has become a standard in benchmarking text generation methods. Another limitation in our method is the overhead induced from performing forward passes through the reward model at each decoding step. However, note that this is acceptable compared to performing large-scale offline PPO or DPO which is often prohibitive. Moreover, this limitation is shared with other RGTG methods.
@@ 6 Conclusion
We have discussed the pitfalls in tokenwise, decoding-time reward-guided text generation (RGTG) with reward models trained on full sequences. These pitfalls can lead to inadequate reward during the autoregressive decoding process and may lead to subpar performance. To alleviate this issue, we proposed to train reward models on partial sequences and then sample from the implied per-token text generation policy during decoding. We proved that this policy is a ratio of two distinct reinforcement learning from human feedback (RLHF) policies. This means that this policy is not equivalent to the standard offline RLHF methods. However, we have also shown that it is intractable to obtain a tokenwise policy that is equivalent to a single RLHF policy. Training a partial-sequence reward model can thus be seen as a tradeoff between avoiding the pitfall of using full-sequence reward model in RGTG and tractability. Our experiment results validated our approach: it performs better than a recent RGTG methods, ARGS, that leverages full-sequence reward models and CD.
[…]
@@
Resources used in this work were provided by the Province of Ontario, the Government of Canada through CIFAR, companies sponsoring the Vector Institute https://vectorinstitute.ai/partners/ and the Natural Sciences and Engineering Council of Canada. AR thanks Apple for support through the Waterloo Apple PhD Fellowship, Natural Sciences and Engineering Council of Canada for its support through the PGS-D program, and the David R. Cheriton Graduate Scholarship. AK thanks Rob Brekelmans for a fruitful discussion. JG thanks Microsoft Research for support through its PhD Scholarship Programme and the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
Appendix A Proofs
See 1
Proof.
See 2
Proof.
In the limit of infinite preference data, maximizing the log-likelihood in (4) is equivalent to minimizing the KL divergence between the learned preference distribution $\sigma$ and the preference data distribution for partial sequences.
$\operatorname*{argmax}_{\phi}\mathop{\mathbb{E}}_{{\mathbf{x}},{\mathbf{y}}_{1},{\mathbf{y}}_{2}\sim P_{\textnormal{data}}}\log\sigma(r_{\phi}({\mathbf{y}}_{1}^{1:i}|{\mathbf{x}})-r_{\phi}({\mathbf{y}}_{2}^{1:j}|{\mathbf{x}}))$ (9)
$=$ $\operatorname*{argmin}_{\phi}-\mathop{\mathbb{E}}_{{\mathbf{x}},{\mathbf{y}}_{1},{\mathbf{y}}_{2}\sim P_{\textnormal{data}}}\log\sigma(r_{\phi}({\mathbf{y}}_{1}^{1:i}|{\mathbf{x}})-r_{\phi}({\mathbf{y}}_{2}^{1:j}|{\mathbf{x}}))$ (10)
$=$ $\operatorname*{argmin}_{\phi}\mathop{\mathbb{E}}_{{\mathbf{x}},{\mathbf{y}}_{1},{\mathbf{y}}_{2}\sim P_{\textnormal{data}}}\log\frac{P_{\textnormal{data}}({\mathbf{y}}_{1}\succcurlyeq{\mathbf{y}}_{2}|{\mathbf{x}},{\mathbf{y}}_{1}^{1:i},{\mathbf{y}}_{2}^{1:j})}{\sigma(r_{\phi}({\mathbf{y}}_{1}^{1:i}|{\mathbf{x}})-r_{\phi}({\mathbf{y}}_{2}^{1:j}|{\mathbf{x}}))}$ (11)
$=$ $\operatorname*{argmin}_{\phi}KL(P_{\textnormal{data}}({\mathbf{y}}_{1}\succcurlyeq{\mathbf{y}}_{2}|{\mathbf{x}},{\mathbf{y}}_{1}^{1:i},{\mathbf{y}}_{2}^{1:j})||\sigma(r_{\phi}({\mathbf{y}}_{1}^{1:i}|{\mathbf{x}})-r_{\phi}({\mathbf{y}}_{2}^{1:j}|{\mathbf{x}})))$ (12)
With a sufficiently expressive reward model, the KL divergence will be zero, and therefore, the distribution $\sigma$ equals the preference data distribution.
$\sigma(r_{\phi}({\mathbf{y}}_{1}^{1:i}|{\mathbf{x}})-r_{\phi}({\mathbf{y}}_{2}^{1:j}|{\mathbf{x}}))=P_{\textnormal{data}}({\mathbf{y}}_{1}\succcurlyeq{\mathbf{y}}_{2}|{\mathbf{x}},{\mathbf{y}}_{1}^{1:i},{\mathbf{y}}_{2}^{1:j})$ (13)
See 3
Proof.
Then, for $i<j$, we have by marginalization:
$=\pi_{\text{ref}}({\mathbf{y}}^{1:i}|{\mathbf{x}})\exp(\beta r({\mathbf{y}}^{1:i}|{\mathbf{x}}))\sum_{{\mathbf{y}}^{i+1:j}}\pi_{\text{ref}}({\mathbf{y}}^{i+1:j}|{\mathbf{x}},{\mathbf{y}}^{1:i})\frac{\exp(\beta r({\mathbf{y}}^{1:j}|{\mathbf{x}}))}{\exp(\beta r({\mathbf{y}}^{1:i}|{\mathbf{x}}))}$
∝̸$\not\propto\pi_{\text{RLHF},i}({\mathbf{y}}^{1:i}|{\mathbf{x}}).∝$
$=\frac{\pi_{\text{ref}}({\mathbf{y}}^{1:i}|{\mathbf{x}})\exp(\beta r({\mathbf{y}}^{1:i}|{\mathbf{x}}))}{\pi_{\text{ref}}({\mathbf{y}}^{1:i-1}|{\mathbf{x}})\exp(\beta r({\mathbf{y}}^{1:i-1}|{\mathbf{x}}))}$ (conditional distribution definition)
This completes the proof of the theorem. ∎
Appendix B Training Details